# How Effective are Large Time Series Models in Hydrology? A Study on Water Level Forecasting in Everglades

### v1

@@ 3.1 Experimental Setup
To ensure consistent evaluation across different models, we considered all variables from the previous 100 days as input and trained all task-specific models to forecast water levels at 5 stations (NP205, P33, G620, NESRS1, and NESRS2, as shown in Figure 2) for the next 7, 14, 21, and 28 days. Time Series Foundation Models are with the same setting in zero-shot inference (without training).
@@
We used Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to evaluate the overall performance of these models. MAE measures the predictive robustness of an algorithm but is relatively insensitive to outliers, whereas RMSE can amplify errors. Additionally, we also include the Symmetric Extremal Dependence Index (SEDI) as a specialized metric for extreme values. It classifies each prediction as either “extreme” or “normal” based on upper or lower quantile thresholds. SEDI values belong to the range $[0,1]$, where higher values indicate better accuracy in identifying extreme water levels.
@@
where $y$ and $\hat{y}$ denote the ground truth and prediction, respectively. The tests, $\hat{y}<y_{\text{up}}^{p}$ and $y<y_{\text{up}}^{p}$ judge whether the predicted and observed values are extremes based on the threshold value $y_{\text{up}}^{p}$. Following the work Han et al. (2024a), we set $10\%$ and $90\%$ as the low and high thresholds, respectively. $N$ is the number of samples.
@@ 4 Results
Surprisingly, the time series foundation model, Chronos, consistently outperforms other models over various lead times across all stations, showcasing exceptional performance in zero-shot inference. Temporally, while the performance of all models declines as the lead time increases, Chronos demonstrates a comparatively slower drop. Spatially, we observe the performance of all models varies across stations with the worst scenarios for the NP205 station, where Chronos still achieves the lowest prediction errors. A possible reason for the lower performance at the NP205 station is its weaker correlation with other water stations (see Figure 8 in Appendix D). Note that the remaining foundation models show relatively poor performance. TimeGPT achieves results comparable to the best task-specific models, NBEATS and PatchTST, while TimesFM, Timer and Moirai perform much worse. Appendix E includes the qualitative analysis to provide a clear observation.
@@
Following the work for extreme weather prediction Han et al. (2024a), we use the SEDI metric to evaluate model performance under extreme values and report the results in Table 2. We observed that Chronos significantly outperforms other models. Moreover, even though TSMixerx shows average performance in overall water level prediction (see Table 1), it performs notably well for predicting extreme water levels at the P33, G620, and NESRS1 stations. To better understand this observation, we visualize the ground truth and model predictions with a lead time of 28 days (see Figure 6 in Appendix B). The visualizations reveal that TSMixerx generally follows the global trend but fails to capture the temporal details (e.g., abrupt changes around time points 50, 130, 150). Notably, despite producing several false alarms between time 140 and 170, TSMixerx successfully predicts most extreme values highlighted in red circles.
@@
While one could attribute the superior performance of Chronos to its tokenization via scaling and quantization or the use of cross entropy in training, a simpler explanation could be that it was exposed to training data that correlates with data from the Everglades National Park. Even though there are no actual Everglades data included, we found that multiple weather datasets were used for its pre-training Ansari et al. (2024). One of them contains daily measurements of precipitation, snow, snow depth, and temperature from climate stations located in 48 states in the USA. However, the other four foundation models in this study were not pre-trained on this dataset or similar datasets.
@@
One key advantage of foundation models is their ability to generate predictions for inputs of varying lengths without being retrained. This differs from task-specific models, which require retraining once the input lengths change. To analyze the impact of input length on prediction performance, we performed experiments predicting 28 days ahead with varying input lengths and reported the results in Figure 5. It shows that the prediction accuracy depends on the input lengths. The MAE values drop as the input length increases from 25 to 100 days, after which the performance stabilizes.
Figure 5: Performance (lead time: 28 days) vs Input Length. We compute the MAE values of ten samples for each input length.
@@ 5 Related Work
Physics-based methods have long been a cornerstone in hydrologic and hydraulics research. For example, 1D and 2D HEC-RAS models have been used to predict water levels Muñoz et al. (2022) and to simulate sewer and street flows in urban systems Huang and Lee (2023). The South Florida Water Management Model (SFWMM) serves as a regional hydrological model designed to simulate water levels in the Everglades SFWMD (2005). Furthermore, researchers have also combined physics-based approaches with ML techniques, such as integrating a hydrologic-hydrodynamic coupling (H2C) model with long short-term memory (LSTM) networks to achieve highly accurate water level predictions in tidal reaches Fei et al. (2023). Despite their utility, these mechanistic models require complex variable integration and rely on predefined assumptions.
@@
The Everglades Forecasting application (EverForecast) Pearlstine et al. (2020) is a statistical model designed to generate spatially continuous simulations of the Everglades water levels in South Florida. Autoregressive Integrated Moving Average (ARIMA) model has been used for water level and discharge predictions Agaj et al. (2024), while the Seasonal Autoregressive Integrated Moving Average (SARIMA) model Azad et al. (2022) was explored for forecasting Malaysia’s mean sea level at coastal stations. In Sekban et al. (2022), the study focuses on predicting monthly dam water levels by comparing various statistical time series models. These models excel at identifying patterns from historical data but are limited in capturing complex nonlinear interactions and dependencies among variables.
@@
ML models such as support vector machines (SVM) Khan and Coulibaly (2006), and extreme gradient boosting (XGBoost) Nguyen et al. (2021) have demonstrated the ability to model nonlinear relationships in water level datasets. We have also experimented with these two ML models, but the results were relatively poor. Besides, the advent of deep learning has revolutionized hydrological forecasting, learning complex spatiotemporal patterns from large datasets. Artificial neural networks (ANNs) Abiy et al. (2022), convolutional neural networks (CNNs) Bassah et al. (2025), recurrent neural networks (RNNs) Shi et al. (2023b), Transformer models Hou et al. (2024), and many hybrid models Shi et al. (2023a) have been used to capture temporal dependencies in hydrological data. Furthermore, Transformer-based models, such as Informer Zhou et al. (2021) and PatchTST Nie et al. (2022), are proposed for time series modeling across various domains. NLinear and DLinear are two models with only linear layers Zeng et al. (2023). Neural Basis Expansion Analysis for Interpretable Time Series Forecasting (NBEATS) Oreshkin et al. (2019) and TimeMixer Chen et al. (2023) are two MLP-based architectures for multi-scale time series modeling. Leveraging large language models (LLMs) for time series forecasting was investigated in Jin et al. (2024), where a time series was treated as a sequence of tokens (values) and pre-trained LLMs were fine-tuned with well-designed prompts Jin et al. (2023).
@@
TimeGPT Garza and Mergenthaler (2023), a pioneering foundation model, employs a transformer-based architecture and was pre-trained on massive time series datasets, supporting both fine-tuning and zero-shot inference. TimesFM Das et al. (2023) follows a similar approach, utilizing a decoder-based architecture to improve time series forecasting. Moment Goswami et al. (2024) introduces a multi-scale learning mechanism, enhancing its ability to model heterogeneous temporal dynamics. Timer Liu et al. (2024a) leverages a hybrid structure incorporating state-space models to improve long-term forecasting stability. Moirai Woo et al. (2024) unifies multiple time series forecasting paradigms within a single foundation model framework. Finally, Chronos Ansari et al. (2024) introduces an innovative design that is optimized for zero-shot forecasting across numerous time series applications.
@@ 6 Discussion
We evaluated 12 task-specific deep learning models and 5 pre-trained time series foundation models for a practical hydrological application, i.e., water level prediction in the Everglades. For task-specific models, performance varies based on architecture and training. They require careful retraining whenever datasets or experimental settings change. For those foundation models, not all of them demonstrate strong performance. Chronos with the zero-shot inference achieves the best performance while others perform poorly. In addition to model architectures, a potential explanation is that Chronos has included training data that highly correlates with data from the Everglades while the training data sets for other foundation models have a low correlation with data from the Everglades or provide limited representations of the hydrological applications. Therefore, incorporating comprehensive data sets for pre-training is crucial to enhance the performance of foundation models. Furthermore, all models struggle to accurately predict extreme high and low values especially when abrupt changes appear. This limitation suggests that the models may have difficulty capturing sudden shifts in the underlying dynamics, potentially due to insufficient training samples for extreme events or inherent model biases. Future improvements could involve incorporating additional physical constraints Yin et al. (2023), leveraging ensemble methods Wu et al. (2023), or integrating uncertainty quantification Shi et al. (2024) to enhance robustness in extreme conditions. Lastly, model performance varies across various locations, demanding context-aware, region-specific adaptions to improve prediction performance Li et al. (2024). These observations further highlight the critical role of domain knowledge in establishing essential physical constraints for more reliable and accurate predictions.
@@ 7 Conclusions
In conclusion, our findings highlight the potential applications of time series foundation models in hydrology. The strong performance of Chronos suggests that certain pre-trained models can leverage their learned representations effectively, even in specialized applications. Moreover, the flexibility of zero-shot inference across different input lengths underscores the high adaptability of foundation models, making them attractive for real-world deployment where data availability and sequence lengths vary. However, further studies are needed to fully harness the potential of these foundation models for hydrologic applications. Exploring strategies to bridge the gap between general-purpose pre-training and the specific requirements of environmental forecasting presents a promising direction for future research.
@@
ML and Everglades researchers were involved in the multidisciplinary collaborations. The domain experts helped in area selection, data acquisition, data preprocessing, and analysis of the results. This work helps domain researchers select the best AI models for water level forecasting in the Everglades and provides practical insights for government agencies interested in using AI-driven hydrological applications.
@@ A.1 Data summary
Rainfall Daily $inches$ 2 NP-205, P33
Potential Evapotranspiration Daily $mm$ 2 NP-205, P33
Pump flows Daily $cfs$ 8 S356, S332B, S332BN, S332C, S332DX1, S332D, S200, S199
Gate flows Daily $cfs$ 11 S343A, S343B, S12A, S12B, S12C, S12D, S [^3] N, S333, S334, S344, S18C
Water levels Daily $ft$ 14 S12A, S12B, S12C, S12D, S333, S334, NESRS1, NESRS2, NP-205, P33, G620, EVER4, TSH, NP62
@@ Appendix F Implementation Details
TimeGPT 100 $1e^{-4}$ Adam Nixtla API [^2] https://docs.nixtla.io/docs/getting-started-about_timegpt Transformer No zero-shot

### v2

@@ 3.1 Experimental Setup
To ensure consistent evaluation across different models, we considered all variables from the previous 100 days as input and trained all task-specific models to forecast water levels at 5 stations (NP205, P33, G620, NESRS1, and NESRS2, as shown in Figure 2) for the next 7, 14, 21, and 28 days. Time Series Foundation Models are with the same setting in zero-shot inference.
@@
Following the work (Shi et al. 2023b, 2025), we used Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to evaluate the overall performance of these models. MAE measures the predictive robustness of an algorithm but is relatively insensitive to outliers, whereas RMSE can amplify errors. Additionally, we also include the Symmetric Extremal Dependence Index (SEDI) as a specialized metric for extreme values (Han et al. 2024a; Zheng et al. 2025). It classifies each prediction as either “extreme” or “normal” based on upper or lower quantile thresholds. SEDI values belong to the range $[0,1]$, where higher values indicate better accuracy in identifying extreme water levels.
@@
where $y$ and $\hat{y}$ denote the ground truth and prediction, respectively. The tests, $\hat{y}<y_{\text{up}}^{p}$ and $y<y_{\text{up}}^{p}$ judge whether the predicted and observed values are extremes based on the threshold value $y_{\text{up}}^{p}$. Following the work (Han et al. 2024b), we set $10\%$ and $90\%$ as the low and high thresholds, respectively. $N$ is the number of samples.
@@ 4 Results
Surprisingly, the time series foundation model, Chronos, consistently outperforms other models over various lead times across all stations, showcasing exceptional performance in zero-shot inference. Temporally, while the performance of all models declines as the lead time increases, Chronos demonstrates a comparatively slower drop. Spatially, we observe the performance of all models varies across stations with the worst scenarios for the NP205 station, where Chronos still achieves the lowest prediction errors. A possible reason for the lower performance at the NP205 station is its weaker correlation with other water stations (see Figure 8 in Appendix D). Note that the remaining foundation models show relatively poor performance. TimeGPT achieves results comparable to the best task-specific models, NBEATS and PatchTST, while TimesFM, Timer and Moirai perform much worse. Due to the limited space, we provide the visualization and qualitative analysis in Appendix E.
@@
Following the work for extreme weather prediction (Han et al. 2024b), we use the SEDI metric to evaluate model performance under extreme values and report the results in Table 2. We observed that Chronos significantly outperforms other models. Moreover, even though TSMixerx shows average performance in overall water level prediction (see Table 1), it performs notably well for predicting extreme water levels at the P33, G620, and NESRS1 stations. To better understand this observation, we visualize the ground truth and model predictions with a lead time of 28 days (see Figure 6 in Appendix B). The visualizations reveal that TSMixerx generally follows the global trend but fails to capture the temporal details (e.g., abrupt changes around time points 50, 130, 150). Notably, despite producing several false alarms between time 140 and 170, TSMixerx successfully predicts most extreme values highlighted in red circles.
@@
While one could attribute the superior performance of Chronos to its tokenization via scaling and quantization or the use of cross entropy in training, a simpler explanation could be that it was exposed to training data that correlates with data from the Everglades National Park. Even though there are no actual Everglades data included, we found that multiple weather datasets were used for its pre-training (Ansari et al. 2024). One of them contains daily measurements of precipitation, snow, snow depth, and temperature from climate stations located in 48 states in the USA. However, the other four foundation models in this study were not pre-trained on this dataset or similar datasets.
@@
One key advantage of foundation models is their ability to generate predictions for inputs of varying lengths without being retrained. To analyze the impact of input length on prediction performance, we performed experiments predicting 28 days ahead with varying input lengths and reported the results in Figure 5. It shows that the prediction accuracy depends on the input lengths. The MAE values drop as the input length increases from 25 to 100 days, after which the performance stabilizes.
Figure 5: Performance (lead time: 28 days) vs Input Length. We computed MAEs of ten samples for each input length.
@@ 5 Related Work
Physics-based methods have long been a cornerstone in hydrologic and hydraulics research. For example, 1D and 2D HEC-RAS models have been used to predict water levels (Muñoz et al. 2022) and to simulate sewer and street flows in urban systems (Huang and Lee 2023). The South Florida Water Management Model (SFWMM) serves as a regional hydrological model designed to simulate water levels in the Everglades (SFWMD 2005). Furthermore, researchers have also combined physics-based approaches with ML techniques, such as integrating a hydrologic-hydrodynamic coupling (H2C) model with long short-term memory (LSTM) networks to achieve highly accurate water level predictions in tidal reaches (Fei, Du, and Gao 2023). Despite their utility, these mechanistic models require complex variable integration and rely on predefined assumptions.
@@
The Everglades Forecasting application (EverForecast) (Pearlstine et al. 2020) is a statistical model designed to generate spatially continuous simulations of the Everglades water levels in South Florida. Autoregressive Integrated Moving Average (ARIMA) model has been used for water level and discharge predictions (Agaj et al. 2024), while the Seasonal Autoregressive Integrated Moving Average (SARIMA) model (Azad et al. 2022) was explored for forecasting Malaysia’s mean sea level at coastal stations. The study (Sekban et al. 2022) focuses on predicting monthly dam water levels by comparing various statistical time series models. These models excel at identifying simple trends from historical data but are limited in capturing complex dynamics among multivariate time series.
@@
ML models such as support vector machines (SVM) (Khan and Coulibaly 2006), and extreme gradient boosting (XGBoost) (Nguyen et al. 2021) have demonstrated the ability to model nonlinear relationships in water level datasets. We have also experimented with these two ML models, but the results were relatively poor. Besides, the advent of deep learning has revolutionized hydrological forecasting, learning complex spatiotemporal patterns from large datasets. Artificial neural networks (ANNs) (Abiy et al. 2022), convolutional neural networks (CNNs) (Bassah et al. 2025), recurrent neural networks (RNNs) (Shi et al. 2023a), Transformer models (Hou et al. 2024), and many hybrid models have been used to capture temporal dependencies in hydrological data. Furthermore, Transformer-based models, such as Informer (Zhou et al. 2021) and PatchTST (Nie et al. 2022), are proposed for time series modeling across various domains. NLinear and DLinear are two models with only linear layers (Zeng et al. 2023). Neural Basis Expansion Analysis for Interpretable Time Series Forecasting (NBEATS) (Oreshkin et al. 2019) and TimeMixer (Chen et al. 2023) are two MLP-based architectures for multi-scale time series modeling. Leveraging large language models (LLMs) for time series forecasting was investigated in (Jin et al. 2024), where a time series was treated as a sequence of tokens (values) and pre-trained LLMs were fine-tuned with well-designed prompts (Jin et al. 2023).
@@
TimeGPT (Garza and Mergenthaler 2023), a pioneering foundation model, employs a transformer-based architecture and was pre-trained on massive time series datasets, supporting both fine-tuning and zero-shot inference. TimesFM (Das et al. 2023) follows a similar approach, utilizing a decoder-based architecture to improve time series forecasting. Moment (Goswami et al. 2024) introduces a multi-scale learning mechanism, enhancing its ability to model heterogeneous temporal dynamics. Timer (Liu et al. 2024a) leverages a hybrid structure incorporating state-space models to improve long-term forecasting stability. Moirai (Woo et al. 2024) unifies multiple time series forecasting paradigms within a single foundation model framework. Finally, Chronos (Ansari et al. 2024) introduces an innovative design that is optimized for zero-shot forecasting across numerous time series applications.
@@ 6 Discussion
We evaluated 12 task-specific DL models and 5 pre-trained foundation models for a hydrological application, i.e., water level prediction in the Everglades. We summarized several takeaways. First, the performance of task-specific models varies based on architecture and training. They require careful retraining whenever datasets or experimental settings change. Second, not all of those foundation models demonstrate strong performance. Chronos with the zero-shot inference achieves the best performance, while others perform poorly. In addition to model architectures, a potential explanation is that Chronos has included training data that highly correlates with data from the Everglades, while the training data sets for other foundation models have a low correlation with data from the Everglades. Third, all models struggle to predict extremely high and low values accurately, suggesting that they may have difficulty capturing abrupt changes or sudden shifts in the underlying dynamics, potentially due to insufficient training samples for extreme events or inherent model biases. Future improvements could involve incorporating additional physical constraints (Yin et al. 2023, 2025), leveraging ensemble methods (Wu et al. 2023), or integrating uncertainty quantification (Shi et al. 2024) to enhance robustness in extreme conditions. Lastly, model performance varies across various locations, demanding context-aware, region-specific adaption to improve prediction performance (Li et al. 2024).
@@ 7 Conclusions
Our findings highlight the potential applications of time series foundation models in hydrology. The strong performance of Chronos suggests that certain pre-trained models can leverage their learned representations effectively, even in specialized applications. Moreover, the flexibility of zero-shot inference across different input lengths underscores the high adaptability of foundation models, making them attractive for real-world deployment where data availability and sequence lengths vary. However, further studies are needed to fully harness the potential of these foundation models for hydrologic applications. Exploring strategies to bridge the gap between general-purpose pre-training and the specific requirements of environmental forecasting presents a promising direction for future research.
@@
ML and Everglades researchers were involved in the multidisciplinary collaborations. Domain experts helped in area selection, data acquisition, data preprocessing, and analysis of the results. This work also helps domain researchers select the best AI models for water level forecasting in the Everglades and provides practical insights for government agencies interested in AI-driven hydrological applications, e.g., flood prediction and management.
@@ A.1 Data summary
Rainfall Daily inchesinches 2 NP-205, P33
Potential Evapotranspiration Daily mmmm 2 NP-205, P33
Pump flows Daily cfscfs 8 S356, S332B, S332BN, S332C, S332DX1, S332D, S200, S199
Gate flows Daily cfscfs 11 S343A, S343B, S12A, S12B, S12C, S12D, S [^3] N, S333, S334, S344, S18C
Water levels Daily ftft 14 S12A, S12B, S12C, S12D, S333, S334, NESRS1, NESRS2, NP-205, P33, G620, EVER4, TSH, NP62
@@ Appendix F Implementation Details
TimeGPT 100 $1e^{-4}$ Adam Nixtla API [^2] https://docs.nixtla.io/docs/getting-started-about˙timegpt Transformer No zero-shot